In an ideal transformer , the no. of turns in primary & secondary are 200 & 1000 respectively .If power input to primary is 10kW at 200V.Calculate


i)output voltage 


ii)current in primary

Grade:11

1 Answers

Yogita Bang
11 years ago

NP/NS = VP/VS

200/1000 = 200/VP   ; Vp = 1000V

 

P=IV; 10000=200*I

I = 50 A

 

If u like the answer then please click YES below